# DESI EMISSION LINE GALAXIES: CLUSTERING DEPENDENCE ON STELLAR MASS AND [OII] LUMINOSITY

### v2

@@ Abstract
We measure the projected two-point correlation functions of emission line galaxies (ELGs) from the Dark Energy Spectroscopic Instrument (DESI) One-Percent Survey and model their dependence on stellar mass and [OII] luminosity. We select $\sim$180,000 ELGs with redshifts of $1.60.8<z<$ 1.6 and define 27 samples according to cuts in redshift and both galaxy properties. Following a framework that describes the conditional [OII] luminosity-stellar mass distribution as a function of halo mass, we simultaneously model the clustering measurements of all samples at fixed redshift. Based on the modeling result, most ELGs in our samples are classified as central galaxies, residing in halos of a narrow mass range with a typical median of $\sim$1012.2-12.4 $h^{-1}M_{\odot}$. We observe a weak dependence of clustering amplitude on stellar mass, which is reflected in the model constraints and is likely a consequence of the 0.5-dex measurement uncertainty in the stellar mass estimates. The model shows a trend between galaxy bias and [OII] luminosity at high redshift ($1.61.2<z<$ 1.6) that is otherwise absent at lower redshifts.
@@ 1 Introduction
In our understanding of cosmology, dark matter dominates the total matter distribution. Primordial fluctuations in the early Universe result in aggregation of dark matter that collapses into dense regions known as halos. Through the cooling and condensation of gas, galaxies form and evolve within these dark matter halos over cosmic time (White & Rees, 1978). While the large-scale structure of dark matter halos has historically been understood analytically (e.g., Press & Schechter 1974; Mo & White 1996; Zentner 2007) and through $N$-body simulations (e.g., Lacey & Cole 1994; Jenkins et al. 2001), baryonic physics complicates the role of galaxies in this paradigm and makes galaxies a biased tracer of the underlying matter distribution. An informative description of this galaxy bias is the relationship between galaxies, their host halos, and the physical properties of both. This relationship remains unclear and an important topic known as the galaxy-halo connection. A deeper understanding of the galaxy-halo connection can provide insights into galaxy formation, evolution, and the nature of dark matter itself (e.g., see the review by Wechsler & Tinker 2018).
@@
One popular route for studying the galaxy-halo connection is through the measurement and modeling of galaxy clustering. Namely, the two-point correlation function (2PCF) quantifies the excess probability (with respect to a random distribution) of observing two galaxies at a given separation (Limber, 1953, 1954; Rubin, 1954). The clustering of galaxies has been measured extensively across spectroscopic surveys and has revealed dependencies on both directly observable and inferred properties, including color, luminosity, stellar mass, and star formation rate (e.g., Zehavi et al. 2011; Mostek et al. 2013; Xu et al. 2018; Clontz et al. 2022). Further, many studies have shown that the clustering of halos themselves is also property-dependent. The mass of the halo is most often considered, but secondary properties such as age, concentration, and formation time are also relevant (e.g., Mo & White 1996; Gao et al. 2005; Wechsler et al. 2006; Xu & Zheng 2018).
@@
Many empirical models have been developed to place galaxies in halos according to these properties and describe the galaxy-halo connection. One such method is known as (sub-halo) abundance matching (SHAM), in which galaxies are ordered according to some property–typically stellar mass–and placed in halos ranked according to halo mass or a similar proxy (Conroy et al., 2006). Other methods, such as the halo occupation distribution (HOD) formalism, describe the probability distribution of a galaxy sample within halos as a function of halo mass (Jing et al., 1998; Seljak, 2000; Berlind & Weinberg, 2002). The halo occupation statistics can then be studied across samples of varying galaxy properties. Conceptually similar, the conditional luminosity function (CLF) framework describes the luminosity distribution of galaxies within halos (Yang et al., 2003). Each of these occupation models, in conjunction with halo clustering statistics, can be used to fit property-dependent galaxy clustering and place constraints on the galaxy-halo connection (Wechsler & Tinker, 2018).
@@
From a galaxy evolution perspective, ELGs are particularly interesting to study. They are actively undergoing star formation, providing a direct probe into this important stage of stellar mass growth. Many models of galaxy formation and evolution have been developed to examine ELGs and their cosmological context (e.g., Gonzalez-Perez et al. 2020).
Newly-formed stars within ELGs can ionize the surrounding gas, producing nebular emission lines that enable an accurate determination of their spectroscopic redshifts. For this reason, and due to their high number density across redshifts of interest, ELGs are being leveraged in modern spectroscopic surveys to study the matter-dominated epoch of the Universe at $z\sim 1$ (Comparat et al., 2013). Specifically, the data from DESI’s survey validation program provide more than 300,000 ELGs to perform our analysis (Raichoor et al., 2023).
The clustering and galaxy-halo connection of ELGs from DESI’s survey validation have been studied individually (Rocher et al., 2023; Ortega-Martinez et al., 2024) and in conjunction with other tracers (Prada et al., 2023; Gao et al., 2023, 2024; Yu et al., 2024; Yuan et al., 2025). Beyond clustering, the intrinsic properties of this sample have been compared to reference star-forming galaxies (Lan et al., 2024). ELG clustering dependence on stellar mass or [OII] luminosity has been studied previously in other surveys (Okumura et al., 2021; Gao et al., 2022); however, no study in DESI has jointly probed the dependence on both properties as presented here. In addition to providing further insight into the galaxy-halo connection, this work will inform the creation of realistic mock catalogs that can be used to calibrate analysis pipelines and test cosmological models.
@@
Throughout this work, we adopt a flat $\Lambda$CDM cosmology consistent with Planck Collaboration et al. (2020), using $\Omega_{m,0}=0.310$, $\Omega_{b,0}=0.049$, $H_{0}=100h$ km Mpc-1 s-1, $h=0.677$, $\sigma_{8}=0.810$, and $n_{s}=0.967$.
@@ 2.1 One-Percent Survey
Upon completion of its five-year operation, the DESI survey will cover $\sim$14,000 deg2 and measure accurate redshifts of nearly 40 million galaxies and quasars (DESI Collaboration et al., 2016a, b). To improve over its predecessors, DESI utilizes 5,000 robotic fiber positioners distributed over ten spectrographs to simultaneously collect 5,000 spectra (Miller et al., 2024; Poppett et al., 2024). The instrument is sensitive from 360 nm to 980 nm and is attached to the 4-meter Mayall telescope at Kitt Peak National Observatory (DESI Collaboration et al., 2022).
@@
The parent sample from which DESI targets are selected is provided by the DESI Legacy Imaging Surveys (Dey et al., 2019), which combine $g$, $r$, and $z$ optical imaging from the Dark Energy Camera (DECam; Flaugher et al. 2015) Legacy Survey (DECaLS), Beijing-Arizona Sky Survey (BASS; Zou et al. 2017), and Mayall $z$-band Legacy Survey (MzLS; Dey et al. 2016) with mid-infrared $W1$, $W2$, $W3$, and $W4$ imaging from the Wide-field Infrared Survey Explorer (WISE; Wright et al. 2010). With this photometry, DESI defines four main extragalactic target classes: the bright galaxy sample (Hahn et al., 2023), luminous red galaxies (Zhou et al., 2023), emission line galaxies (Raichoor et al., 2023), and quasars (Chaussidon et al., 2023). Stars in the Milky Way are also observed (Cooper et al., 2023).
@@
This work utilizes large-scale structure (LSS) catalogs [^3] https://data.desi.lbl.gov/public/edr/vac/edr/lss/v2.0/ from the DESI survey validation program that covers $\sim$140 deg2 of the planned survey area, also known as the One-Percent Survey (DESI Collaboration et al., 2024a). The spectroscopic data were processed by the pipeline described in Guy et al. (2023) and fit by Redrock [^4] https://github.com/desihub/redrock/, the spectral template software, to classify each object and determine redshifts. These data were made publicly available as part of the DESI Early Data Release (DESI Collaboration et al., 2024b) [^5] https://data.desi.lbl.gov/doc/releases/edr/. The footprint consists of 20 regions (“rosettes”) and is shown in Figure 1. In order to balance the redshift coverage of the ELGs with other targets during fiber assignment, each ELG is assigned one of two priority designations, low ($\tt{ELG\_LOP}$) or very low ($\tt{ELG\_VLO}$), in the form of targeting bits (Myers et al., 2023). To perform the clustering measurements, each LSS catalog is accompanied by several “random” catalogs that are created to match the data’s angular coverage and redshift distribution but with an otherwise random spatial distribution (Ross et al., 2024).
As in Rocher et al. (2023), we define an annulus within each rosette (inset of Figure 1) that spans 0.2°$\degree$ to 1.5°$\degree$ radially outward from the rosette center and corresponds to an angular area of $\sim$7 deg2. Within each annulus, $\sim$95of ELG targets had their spectra reliably measured (DESI Collaboration et al., 2024a), resulting in high completeness. To construct our overall sample, we use ELGs from the One-Percent LSS catalogs that have either $\tt{ELG\_LOP}$ or $\tt{ELG\_VLO}$ priority designation and are contained within the annuli defined above.
@@
Figure 1: Footprint of the 20 One-Percent Survey rosettes in blue (Mollweide projection). The galactic plane is shown in green. The inset image shows the angular distribution of spectroscopically-confirmed ELGs within one rosette; the galaxies observed in the highly complete annulus are highlighted in blue.
@@ 2.2 ELG Samples
ELG redshifts can be measured quickly and accurately via strong emission lines and, in particular, the [OII] doublet at rest-frame wavelengths of 3,726 Å and 3,729 Å (Raichoor et al., 2023). Further, ELGs are abundant when the average star-formation rate of the Universe peaked at redshift of $z\sim 2$ (Madau & Dickinson, 2014), making them especially useful out to $z\sim 1.6$ before the [OII] doublet is redshifted out of DESI’s wavelength coverage. The ELG sample targeted by DESI lies between $1.60.6<z<$ 1.6, extending to sources one magnitude fainter and ten times denser than that of its predecessor, the Extended Baryon Oscillation Spectroscopic Survey (eBOSS; Dawson et al. 2016; Raichoor et al. 2017). This surface density is sufficient to reach the sub-percent precision for measurements of the Baryon Acoustic Oscillation feature pursued by DESI (Raichoor et al., 2023). The number density of the ELGs in the DESI One-Percent survey as a function of redshift is shown in Figure 2.
Figure 2: Number density of spectroscopically-confirmed ELGs in the DESI One-Percent Survey. Only galaxies within rosette annuli (Figure 1) are included. Colored regions mark the redshift ranges considered in this work.
@@
We do not use all ELGs from the LSS catalogs, but rather require those with $F_{\text{[OII]}}$ signal-to-noise $>3$ to ensure reasonable precision in the [OII] flux measurement. The median fractional uncertainty in total [OII] flux is $\sim$6%. Galaxies were then divided into three redshift ranges: $1.00.8<z<$ 1.0, $1.21.0<z<$ 1.2, and $1.61.2<z<$ 1.6.
Within each redshift range, the physical properties provided by FastSpecFit were compared with those from COSMOS2020, the latest release of the Cosmic Evolution Survey (Weaver et al., 2022, 2023). This comparison revealed that, relative to COSMOS2020, DESI has a measurement scatter of $\sigma_{\log M_{*}}\sim 0.5$ in stellar mass. Such a scatter can flatten the $\log L_{\rm[OII]}$-$\log M_{*}$ distribution of our samples and influence our model constraints. A detailed discussion is given in Appendix A. Uncertainty in the stellar mass estimates and other selection criteria naturally lead to incompleteness in the DESI selection, meaning this work reflects the DESI sample rather than an intrinsic star-forming galaxy population.
@@
Figure 3: Selection boundaries for the nine samples within each redshift group as a function of $\log L_{\text{[OII]}}$ and $\log M_{*}$. Left, middle, and right panels are at redshift $1.00.8<z<$ 1.0, $1.21.0<z<$ 1.2, and $1.61.2<z<$ 1.6, respectively.
@@
Z1 Parent 0.91 7.63aaExcluding less than three extreme outliers in the FastSpecFit value-added catalog. 12.52 9.87 40.60 43.23 41.86 71,995 11.60
Z1M1L1 0.89 8.50 9.59 9.20 41.52 41.78 41.68 7,448 1.18
Z1M1L2 0.91 8.50 9.59 9.22 41.78 41.95 41.86 7,182 1.15
Z1M1L3 0.92 8.50 9.59 9.26 41.95 42.80 42.11 7,290 1.16
Z1M2L1 0.89 9.59 10.11 9.85 41.52 41.81 41.70 7,561 1.22
Z1M2L2 0.91 9.59 10.11 9.84 41.81 41.99 41.90 7,289 1.19
Z1M2L3 0.93 9.59 10.11 9.84 41.99 42.80 42.15 7,413 1.20
Z1M3L1 0.89 10.11 12.52 10.61 41.52 41.79 41.68 7,166 1.16
Z1M3L2 0.91 10.11 12.52 10.58 41.79 41.98 41.88 7,396 1.20
Z1M3L3 0.93 10.11 12.52 10.61 41.98 42.80 42.14 7,355 1.22
@@
Z2M1L1 1.09 8.50 9.63 9.26 41.85 42.01 41.94 5,604 0.74
Z2M1L2 1.10 8.50 9.63 9.28 42.01 42.15 42.08 5,261 0.70
Z2M1L3 1.11 8.50 9.63 9.31 42.15 42.80 42.29 5,541 0.73
Z2M2L1 1.09 9.63 10.11 9.85 41.85 42.04 41.96 5,617 0.75
Z2M2L2 1.10 9.63 10.11 9.86 42.04 42.19 42.11 5,127 0.69
Z2M2L3 1.11 9.63 10.11 9.87 42.19 42.80 42.33 5,477 0.73
Z2M3L1 1.09 10.11 12.61 10.66 41.85 42.03 41.95 5,568 0.76
Z2M3L2 1.10 10.11 12.61 10.64 42.03 42.19 42.11 5,387 0.72
Z2M3L3 1.11 10.11 12.61 10.65 42.19 42.80 42.34 5,566 0.75
Z3 Parent 1.38 7.82aaExcluding less than three extreme outliers in the FastSpecFit value-added catalog. 12.99 10.04 41.08 43.81 42.26 79,572 4.50
Z3M1L1 1.35 8.50 9.75 9.33 42.08 42.22 42.15 6,804 0.38
Z3M1L2 1.38 8.50 9.75 9.35 42.22 42.37 42.29 7,000 0.39
Z3M1L3 1.41 8.50 9.75 9.38 42.37 42.80 42.50 6,740 0.38
Z3M2L1 1.35 9.75 10.28 9.99 42.08 42.25 42.17 6,913 0.39
Z3M2L2 1.38 9.75 10.28 10.01 42.25 42.40 42.32 6,497 0.37
Z3M2L3 1.41 9.75 10.28 10.02 42.40 42.80 42.53 6,727 0.38
Z3M3L1 1.37 10.28 12.91 10.85 42.08 42.25 42.17 6,762 0.38
Z3M3L2 1.40 10.28 12.91 10.84 42.25 42.42 42.33 7,071 0.40
Z3M3L3 1.44 10.28 12.91 10.84 42.42 42.80 42.55 6,595 0.39
Note. — Horizontal lines separate redshift groups: Z1 $\in(0.8,1.0)$, Z2 $\in(1.0,1.2)$, Z3 $\in(1.2,1.6)$. The number of galaxies per sample is listed as $N_{g}$. The measured number densities, weighted according to Equation (5), are listed as $n_{g}$. Stellar masses, [OII] luminosities, and number densities are in units of $M_{\odot}$, erg s-1, and $10^{-4}$ ($h^{-1}$pc)-3, respectively. Parent samples are defined with the annulus and signal-to-noise requirements imposed (Sections 2.1 and 2.2).
After assigning galaxies to each sample, the number densities $n_{g}(z)$ no longer match that of the random catalogs conructed for the entire sample. To match the $n_{g}(z)$ for each of the nine samples in each redshift bin, objects in the random catalogs were divided into fine redshift bins ($\Delta z=0.01$) and randomly down-sampled.

### v3

# DESI EMISSION-LINE GALAXIES: CLUSTERING DEPENDENCE ON STELLAR MASS AND [OII] LUMINOSITY

@@ Abstract
We measure the projected two-point correlation functions of emission-line galaxies (ELGs) from the Dark Energy Spectroscopic Instrument One-Percent Survey and model their dependence on stellar mass and [OII] luminosity. We select $\sim$180,000 ELGs with redshifts of $0.8<z<1.6$ and define 27 samples according to cuts in redshift and both galaxy properties. Following a framework that describes the conditional [OII] luminosity-stellar mass distribution as a function of halo mass, we simultaneously model the clustering measurements of all samples at fixed redshift. Based on the modeling result, most ELGs in our samples are classified as central galaxies, residing in halos of a narrow mass range with a typical median of $\sim$1012.2-12.4 $h^{-1}M_{\odot}$. We observe a weak dependence of clustering amplitude on stellar mass, which is reflected in the model constraints and is likely a consequence of the 0.5 dex measurement uncertainty in the stellar mass estimates. The model shows a trend between galaxy bias and [OII] luminosity at high redshift ($1.2<z<1.6$) that is otherwise absent at lower redshifts.
@@ 1 Introduction
In our understanding of cosmology, dark matter dominates the total matter distribution. Primordial fluctuations in the early Universe result in aggregation of dark matter that collapses into dense regions known as halos. Through the cooling and condensation of gas, galaxies form and evolve within these dark matter halos over cosmic time (White & Rees, 1978). While the large-scale structure (LSS) of dark matter halos has historically been understood analytically (e.g., Press & Schechter 1974; Mo & White 1996; Zentner 2007) and through $N$-body simulations (e.g., Lacey & Cole 1994; Jenkins et al. 2001), baryonic physics complicates the role of galaxies in this paradigm and makes galaxies a biased tracer of the underlying matter distribution. An informative description of this galaxy bias is the relationship between galaxies, their host halos, and the physical properties of both. This relationship remains unclear and an important topic known as the galaxy-halo connection. A deeper understanding of the galaxy-halo connection can provide insights into galaxy formation, evolution, and the nature of dark matter itself (see, e.g., the review by Wechsler & Tinker 2018).
@@
One popular route for studying the galaxy-halo connection is through the measurement and modeling of galaxy clustering. Namely, the two-point correlation function (2PCF) quantifies the excess probability (with respect to a random distribution) of observing two galaxies at a given separation (Limber, 1953, 1954; Rubin, 1954). The clustering of galaxies has been measured extensively across spectroscopic surveys and has revealed dependencies on both directly observable and inferred properties, including color, luminosity, stellar mass, and star formation rate (e.g., Zehavi et al. 2011; Mostek et al. 2013; Xu et al. 2018; Clontz et al. 2022). Further, many studies have shown that the clustering of halos themselves is also property dependent. The mass of the halo is most often considered, but secondary properties such as age, concentration, and formation time are also relevant (e.g., Mo & White 1996; Gao et al. 2005; Wechsler et al. 2006; Xu & Zheng 2018).
@@
Many empirical models have been developed to place galaxies in halos according to these properties and describe the galaxy-halo connection. One such method is known as (subhalo) abundance matching (SHAM), in which galaxies are ordered according to some property–typically stellar mass–and placed in halos ranked according to halo mass or a similar proxy (Conroy et al., 2006). Other methods, such as the halo occupation distribution (HOD) formalism, describe the probability distribution of a galaxy sample within halos as a function of halo mass (Jing et al., 1998; Seljak, 2000; Berlind & Weinberg, 2002). The halo occupation statistics can then be studied across samples of varying galaxy properties. Conceptually similar, the conditional luminosity function (CLF) framework describes the luminosity distribution of galaxies within halos (Yang et al., 2003). Each of these occupation models, in conjunction with halo clustering statistics, can be used to fit property-dependent galaxy clustering and place constraints on the galaxy-halo connection (Wechsler & Tinker, 2018).
@@
From a galaxy evolution perspective, ELGs are particularly interesting to study. They are actively undergoing star formation, providing a direct probe into this important stage of stellar mass growth. Many models of galaxy formation and evolution have been developed to examine ELGs and their cosmological context (e.g., Gonzalez-Perez et al. 2020). For star-forming galaxies like ELGs, stellar mass and star formation rate are major observable properties that describe their evolutionary status. Inference of the connection between these properties and the host dark matter halos from galaxy clustering data informs the role of halo environment in shaping such properties and provides an empirical test for galaxy formation models.
Newly formed stars within ELGs can ionize the surrounding gas, producing nebular emission lines that enable an accurate determination of their spectroscopic redshifts. For this reason, and due to their high number density across redshifts of interest, ELGs are being leveraged in modern spectroscopic surveys to study the matter-dominated epoch of the Universe at $z\sim 1$ (Comparat et al., 2013). Specifically, the data from DESI’s survey validation program provide more than 300,000 ELGs for our analysis (Raichoor et al., 2023).
The clustering and galaxy-halo connection of ELGs from DESI’s survey validation have been studied individually (Rocher et al., 2023; Ortega-Martinez et al., 2024) and in conjunction with other tracers (Gao et al., 2023, 2024; Prada et al., 2025; Yu et al., 2024; Yuan et al., 2025). Beyond clustering, the intrinsic properties of this sample have been compared to reference star-forming galaxies (Lan et al., 2024). ELG clustering dependence on stellar mass or [OII] luminosity has been studied previously in other surveys (Okumura et al., 2021; Gao et al., 2022); however, no study in DESI has jointly probed the dependence on both properties as presented here. Doing so will provide a more detailed and complete description of the galaxy-halo connection, enabling a deeper understanding of galaxy formation. It will also inform the creation of realistic mock catalogs that can be used to calibrate analysis pipelines and test cosmological models.
@@
[…]
@@ 2.1 One-Percent Survey
Upon completion of its 5 yr operation, the DESI survey will have covered $\sim$14,000 deg2 and measured accurate redshifts for nearly 40 million galaxies and quasars (DESI Collaboration et al., 2016a, b). To improve over its predecessors, DESI utilizes 5000 robotic fiber positioners distributed over 10 spectrographs to simultaneously collect 5000 spectra (Miller et al., 2024; Poppett et al., 2024). The instrument is sensitive from 360 to 980 nm and is attached to the 4 m Mayall telescope at Kitt Peak National Observatory (DESI Collaboration et al., 2022).
@@
The parent sample from which DESI targets are selected is provided by the DESI Legacy Imaging Surveys (Dey et al., 2019), which combine $g$, $r$, and $z$ optical imaging from the Dark Energy Camera (Flaugher et al., 2015) Legacy Survey (or DECaLS), the Beijing-Arizona Sky Survey (or BASS; Zou et al. 2017), and the Mayall $z$-band Legacy Survey (or MzLS; Dey et al. 2016) with mid-infrared $W1$, $W2$, $W3$, and $W4$ imaging from the Wide-field Infrared Survey Explorer (or WISE; Wright et al. 2010). With this photometry, DESI defines four main extragalactic target classes: the bright galaxy sample (Hahn et al., 2023), luminous red galaxies (Zhou et al., 2023), ELGs (Raichoor et al., 2023), and quasars (Chaussidon et al., 2023). Stars in the Milky Way are also observed (Cooper et al., 2023).
@@
This work utilizes LSS catalogs [^3] https://data.desi.lbl.gov/public/edr/vac/edr/lss/v2.0/ from the DESI survey validation program that covers $\sim$140 deg2 of the planned survey area, also known as the One-Percent Survey (DESI Collaboration et al., 2024a). The spectroscopic data were processed by the pipeline described in Guy et al. (2023) and fit by Redrock [^4] https://github.com/desihub/redrock/, the spectral template software, to classify each object and determine redshifts. These data were made publicly available as part of the DESI Early Data Release (DESI Collaboration et al., 2024b) [^5] https://data.desi.lbl.gov/doc/releases/edr/. The footprint consists of 20 regions (“rosettes,” with each being slightly larger than the field of view of the DESI focal plane) and is shown in Figure 1. To balance the redshift coverage of the ELGs with other targets during fiber assignment, each ELG is assigned one of two priority designations, low ($\tt{ELG\_LOP}$) or very low ($\tt{ELG\_VLO}$), in the form of targeting bits (Myers et al., 2023). Because each rosette in the One-Percent survey received a minimum of 11 observational visits, the ELG sample is highly complete despite having lower priority than other target classes (DESI Collaboration et al., 2024a). To perform the clustering measurements, each LSS catalog is accompanied by several “random” catalogs that are created to match the data’s angular coverage (e.g., regions around bright stars are masked) and redshift distribution but with an otherwise random spatial distribution (Ross et al., 2024).
[…]
@@
Figure 1: Footprint of the 20 One-Percent Survey rosettes in blue (Mollweide projection). The Galactic plane is shown in green. The inset image shows the angular distribution of spectroscopically confirmed ELGs within one rosette; the galaxies observed in the highly complete annulus are highlighted in blue.
@@ 2.2 ELG Samples
ELG redshifts can be measured quickly and accurately via strong emission lines, and in particular the [OII] doublet at rest-frame wavelengths of 3726 and 3729 Å (Raichoor et al., 2023). Further, ELGs are abundant at the epoch when the average star formation rate of the Universe peaked, at a redshift of $z\sim 2$ (Madau & Dickinson, 2014), making them especially useful out to $z\sim 1.6$ before the [OII] doublet is redshifted out of DESI’s wavelength coverage. The ELG sample targeted by DESI lies between $0.6<z<1.6$, extending to sources 1 magnitude fainter and 10 times denser than that of its predecessor, the Extended Baryon Oscillation Spectroscopic Survey (or eBOSS; Dawson et al. 2016; Raichoor et al. 2017). This surface density is sufficient to reach the subpercent precision necessary for measurements of the baryon acoustic oscillation feature pursued by DESI (Raichoor et al., 2023). The number density of the ELGs in the DESI One-Percent survey as a function of redshift is shown in Figure 2.
Figure 2: Number density of spectroscopically confirmed ELGs in the DESI One-Percent Survey. Only galaxies within rosette annuli (Figure 1) are included. Colored regions mark the redshift ranges considered in this work.
@@
We do not use all ELGs from the LSS catalogs, but rather require those with a $F_{\text{[OII]}}$ signal-to-noise ratio >3>3 to ensure reasonable precision in the [OII] flux measurement (removing an additional $\sim$0.6% of objects). The median fractional uncertainty in total [OII] flux is $\sim$6%. Galaxies were then divided into three redshift ranges: $0.8<z<1.0$, $1.0<z<1.2$, and $1.2<z<1.6$.
Within each redshift range, the physical properties provided by FastSpecFit were compared with those from COSMOS2020, the latest release of the Cosmic Evolution Survey (Weaver et al., 2022, 2023). This comparison revealed that, relative to COSMOS2020, DESI has a measurement scatter of $\sigma_{\log M_{*}}\sim 0.5$ in stellar mass. Such a scatter can flatten the $\log L_{\rm[OII]}$-$\log M_{*}$ distribution of our samples and influence our model constraints. A detailed discussion is given in Appendix A. Uncertainties in the stellar mass estimates and other selection criteria naturally lead to incompleteness in the DESI selection, meaning this work reflects the DESI sample rather than an intrinsic star-forming galaxy population.
@@
Figure 3: Selection boundaries for the nine samples within each redshift group as a function of $\log L_{\text{[OII]}}$ and $\log M_{*}$. The horizontal error bars illustrate the 0.5 dex measurement scatter in stellar mass as quantified in Appendix A with respect to COSMOS2020. Left, middle, and right panels correspond to redshift ranges $0.8<z<1.0$, $1.0<z<1.2$, and $1.2<z<1.6$, respectively.
@@
Z1 Parent 0.91 7.63aaExcluding fewer than three extreme outliers in the FastSpecFit value-added catalog. 12.52 9.87 40.60 43.23 41.86 71,995 11.60
Z1M1L1 0.89 8.50 9.59 9.20 41.52 41.78 41.68 7448 1.18
Z1M1L2 0.91 8.50 9.59 9.22 41.78 41.95 41.86 7182 1.15
Z1M1L3 0.92 8.50 9.59 9.26 41.95 42.80 42.11 7290 1.16
Z1M2L1 0.89 9.59 10.11 9.85 41.52 41.81 41.70 7561 1.22
Z1M2L2 0.91 9.59 10.11 9.84 41.81 41.99 41.90 7289 1.19
Z1M2L3 0.93 9.59 10.11 9.84 41.99 42.80 42.15 7413 1.20
Z1M3L1 0.89 10.11 12.52 10.61 41.52 41.79 41.68 7166 1.16
Z1M3L2 0.91 10.11 12.52 10.58 41.79 41.98 41.88 7396 1.20
Z1M3L3 0.93 10.11 12.52 10.61 41.98 42.80 42.14 7355 1.22
@@
Z2M1L1 1.09 8.50 9.63 9.26 41.85 42.01 41.94 5604 0.74
Z2M1L2 1.10 8.50 9.63 9.28 42.01 42.15 42.08 5261 0.70
Z2M1L3 1.11 8.50 9.63 9.31 42.15 42.80 42.29 5541 0.73
Z2M2L1 1.09 9.63 10.11 9.85 41.85 42.04 41.96 5617 0.75
Z2M2L2 1.10 9.63 10.11 9.86 42.04 42.19 42.11 5127 0.69
Z2M2L3 1.11 9.63 10.11 9.87 42.19 42.80 42.33 5477 0.73
Z2M3L1 1.09 10.11 12.61 10.66 41.85 42.03 41.95 5568 0.76
Z2M3L2 1.10 10.11 12.61 10.64 42.03 42.19 42.11 5387 0.72
Z2M3L3 1.11 10.11 12.61 10.65 42.19 42.80 42.34 5566 0.75
Z3 Parent 1.38 7.82aaExcluding fewer than three extreme outliers in the FastSpecFit value-added catalog. 12.99 10.04 41.08 43.81 42.26 79,572 4.50
Z3M1L1 1.35 8.50 9.75 9.33 42.08 42.22 42.15 6804 0.38
Z3M1L2 1.38 8.50 9.75 9.35 42.22 42.37 42.29 7000 0.39
Z3M1L3 1.41 8.50 9.75 9.38 42.37 42.80 42.50 6740 0.38
Z3M2L1 1.35 9.75 10.28 9.99 42.08 42.25 42.17 6913 0.39
Z3M2L2 1.38 9.75 10.28 10.01 42.25 42.40 42.32 6497 0.37
Z3M2L3 1.41 9.75 10.28 10.02 42.40 42.80 42.53 6727 0.38
Z3M3L1 1.37 10.28 12.91 10.85 42.08 42.25 42.17 6762 0.38
Z3M3L2 1.40 10.28 12.91 10.84 42.25 42.42 42.33 7071 0.40
Z3M3L3 1.44 10.28 12.91 10.84 42.42 42.80 42.55 6595 0.39
Note. — Horizontal lines separate redshift groups: Z1 $\in(0.8,1.0)$, Z2 $\in(1.0,1.2)$, Z3 $\in(1.2,1.6)$. The number of galaxies per sample is listed as $N_{g}$. The measured number densities, weighted according to Equation (5), are listed as $n_{g}$. Stellar masses, [OII] luminosities, and number densities are in units of solar masses ($M_{\odot}$), erg per second (erg s-1), and $10^{-4}$ ($h^{-1}$Mpc)-3, respectively. Parent samples are defined with the annulus and signal-to-noise requirements imposed (Sections 2.1 and 2.2).
After assigning galaxies to each sample, the number densities $n_{g}(z)$ no longer match those of the random catalogs constructed for the entire sample. To match the $n_{g}(z)$ for each of the nine samples in each redshift bin, objects in the random catalogs were divided into fine redshift bins ($\Delta z=0.01$) and randomly down-sampled.